a rock climber burns 60 calories for every 5 minutes of rock climbing.
write a proportion on the line below to determine how many calories a rock climber would burn in 7 minutes.

proportion________________



now use the proportion you wrote to determine how many calories the rock climber would burn in 7 minutes. write your answer on the line below.
SHOW ALL WORK





answer________________calories

Answer by checkley77(12569)

60/5=x/7
5x=7*60
5x=420
x=420/5
x=84 calories will be burned in 7 minutes.